OpenStudy (anonymous):

Jim bought 2 packs of batteries from a store. The price of each pack was the same. After he bought the batteries, his account balance showed a change of −$23.94. What would have been the change to Jim's account balance had he only bought 1 pack of batteries from the store? −$21.94 −$11.97 $11.97 $21.94

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (anonymous):

divide by 2 and add a negative sign so -11.97

OpenStudy (anonymous):

Which of the following describes a situation in which the total distance a car travels is 0 meters from its starting point? The car first travels a distance of 11 meters and then travels 9 meters in the opposite direction. The car first travels 9 meters backward and then travels another 9 meters in the same direction. The car first travels 11 meters backward and then travels 9 meters in the same direction. The car first travels a distance of 9 meters and then travels 9 meters in the opposite direction. @chrisk123

OpenStudy (anonymous):

the last one

OpenStudy (anonymous):

c

OpenStudy (anonymous):

@chrisk123

OpenStudy (anonymous):

Multiply 1 over 3 multiplied by 1 over 5. negative 1 over 10 negative 1 over 15 1 over 15 1 over 10 Question 10 (Multiple Choice Worth 5 points) (01.01 LC)Look at the following numbers: −4, −5, 0, 5 Which pair of numbers has a sum of zero? 5, −5 −5, 0 −4, 5 −4, 0 Question 11 (Multiple Choice Worth 5 points) (01.02 LC)Simplify negative 4 and 1 over 4 − negative 9 and 1 over 2. negative 13 and 1 over 3 negative 5 and 1 over 4 13 and 1 over 3 5 and 1 over 4 Question 12 (Multiple Choice Worth 5 points) (01.01 HC)Point N on the number line shows Rob's score after the first round of a quiz: A number line is shown from negative 8 to 0 to positive 8. There are increments of 1 on either side of the number line. The even numbers are labeled on either side of the number line. Point N is shown on 1. In round two, he lost 6 points. Which expression shows how many total points he has at the end of round two? 1 + (−6) = −5, because −5 is 6 units to the left of 1 1 + 6 = −5, because −5 is 6 units to the left of 1 1 + (−5) = −6, because −6 is 5 units to the left of 1 1 + 5 = −6, because −6 is 5 units to the left of 1

OpenStudy (anonymous):

@chrisk123 sorry thts it

OpenStudy (anonymous):

oint N on the number line shows Rob's score after the first round of a quiz: A number line is shown from negative 8 to 0 to positive 8. There are increments of 1 on either side of the number line. The even numbers are labeled on either side of the number line. Point N is shown on 1. In round two, he lost 6 points. Which expression shows how many total points he has at the end of round two? 1 + (−6) = −5, because −5 is 6 units to the left of 1 1 + 6 = −5, because −5 is 6 units to the left of 1 1 + (−5) = −6, because −6 is 5 units to the left of 1 1 + 5 = −6, because −6 is 5 units to the left of 1

OpenStudy (anonymous):

is there a pic